# Learn from Global Correlations: Enhancing Evolutionary Algorithm via Spectral GNN

### v3

@@ 1 Introduction
Graph Neural Networks (GNNs) have been widely applied in fields such as social network analysis, recommendation systems, and protein structure prediction because of their capacity to effectively model complex relational structures using graph representations[1, 2, 3, 4, 5]. In GNNs, dependencies between nodes are captured by propagating and aggregating information across the network, thereby enabling adaptive learning and optimization based on node interactions. Similarly, biological evolution involves genetic interactions characterized by variation, recombination, and natural selection, which are processes that collectively facilitate the evolutionary optimization of genotypes toward greater fitness[6, 7, 8]. Given these parallels, we propose that GNNs can be understood through an evolutionary perspective, where the propagation of information among nodes mirrors the diversity-exploring mechanisms of crossover and mutation, as well as the adaptive optimization mechanism of directional selection in evolution.
@@
Evolutionary adaptation thrives on two complementary strategies: exploration through genetic diversity (e.g., avian wing polymorphism spanning short, medium, and long morphotypes) and exploitation via trait optimization[9, 10, 11].Consider three avian archetypes: short-winged penguins excel in aquatic foraging, medium-winged gulls balance aerial maneuverability with diving efficiency, and long-winged albatrosses dominate transoceanic flight—each adaptation emerging from mutation-driven exploration and sustained through environmental selection. Similarly, giraffes’ neck elongation, refined iteratively over generations, exemplifies exploitation’s role in amplifying fitness-critical traits. Evolutionary algorithms (EAs) formalize this duality: crossover/mutation operators probe global solution spaces (exploration), while selection pressures refine local optima (exploitation)[12, 13, 14].
@@
Conventional EAs fail to fully exploit the global relationships among individuals in a population, relying on limited information from a few individuals for updating a single solution[15, 10, 16], thus hindering comprehensive learning. Additionally, existing EAs optimize functions in a black-box manner, with parameter selection relying on extensive experimental results. Inappropriate parameter choices can easily disrupt the balance between exploration and exploitation, potentially leading to suboptimal results [17, 18, 19].To address these limitations, we reinterpret evolutionary algorithms through a GNN lens, proposing Graph Neural Evolution (GNE). In GNE, each candidate solution in population is represented as a node in a graph, with the corresponding value in the adjacency matrix between two nodes reflecting the similarity of them. This approach fully leverages global correlations among individuals within the population, overcoming the constraints of local updates. These global correlations resemble biological reproductive isolation, where genetically similar individuals preferentially exchange traits[20]. Moreover, we compute the normalized Laplacian matrix from the population’s adjacency matrix and decompose it into frequency components through its eigenvalues and eigenvectors in the spectral domain. Similar to spectral GNNs, GNE balances high-frequency exploration (diverse traits) and low-frequency exploitation (consistent traits) via filter functions. Notably, this explicit spectral aggregation strategy can control global-scale features through frequency components, making the exploration–exploitation control both interpretable and effective.
@@
This spectral perspective provides a unified mathematical framework: the balance between exploration and exploitation in evolutionary algorithms is equivalent to designing an appropriate filter that assigns suitable weights to the different frequency components of the population, where strengthening the high-frequency components corresponds to reinforcing exploration behavior, and strengthening the low-frequency components corresponds to reinforcing exploitation behavior. By formalizing this duality, GNE provides interpretable control over the exploration-exploitation balance while benefiting from global correlations. Through rigorous evaluations on nine benchmark functions (e.g., Sphere, Rastrigin, and Rosenbrock), GNE consistently outperforms classical algorithms, including Genetic Algorithm (GA) [12], Differential Evolution (DE) [14], Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [21], and advanced algorithms such as Evolution Strategy Based on Search Direction Adaptation (SDAES) [22] and Learning Adaptive Differential Evolution by Natural Evolution Strategies (RL-SHADE) [23] across scenarios with original, noise-corrupted, and deviated-optimum conditions. GNE leads other algorithms by multiple orders of magnitude in solution quality (e.g., 3.07e-20 vs. 1.51e-07 mean on Sphere). This framework is broadly applicable to complex optimization problems in the real world, such as engineering design, and optimization of parameters and architecture of neural networks[24, 25, 26].To summarize, our contributions are threefold:
@@ 2 Related Works
Evolutionary Algorithms. EAs are a family of metaheuristic optimization techniques inspired by natural evolutionary processes, incorporating operators such as natural selection and mutation. GA is one of the earliest and most representative EAs, and is particularly effective in solving discrete and multimodal optimization problems[12]. In recent years, DE[14, 16, 27] and Evolution Strategy (ES)[28, 29, 30] have emerged as prominent research domains. DE generates candidate solutions through differential mutation and selects the fittest among them. Numerous variants have been developed, including Adaptive Differential Evolution with Optional External Archive (JADE) [31], which incorporates external archives and adaptive parameter updates; Success-History Based Parameter Adaptation for Differential Evolution (SHADE) [32], which utilizes historically successful parameter settings to refine control parameters; and RL-SHADE [23], which further enhances DE performance by integrating natural evolution strategies for more effective parameter optimization. Among ES methods, Covariance Matrix Adaptation Evolution Strategy (CMA-ES) is one of the most widely recognized algorithms [21], adjusting the covariance matrix to effectively explore the solution space. Several variants have been developed, including OpenES for large-scale parallel computation [33]; Parameter-Exploring Policy Gradients (PEPG) for estimating gradients in the parameter space [34];Mixture Model based Evolution Strategy (MMES) [35], which accelerates CMA-ES in large-scale optimization tasks; and SDAES [22], which reduces both time and space complexity in high-dimensional settings. From GA to DE, ES, and their variants, the design of evolutionary algorithms has consistently centered on balancing exploration and exploitation. Exploration introduces diversity to expand the search space and prevent premature convergence to local optima, while exploitation refines promising solutions to gradually approach the global optimum[9].
@@
Graph Neural Networks. GNNs process graph-structured data from two principal perspectives: spatial and spectral[36, 37, 38]. Spatial GNNs are more widely known, with the Graph Convolutional Network (GCN) [1] being a representative example, which aggregates information from neighboring nodes to learn local structures. The Graph Attention Network (GAT) [2] further enhances GCNs by incorporating self-attention mechanisms, while GraphSAGE generates node embeddings by sampling and aggregating features from local neighborhoods, thereby enabling generalization to previously unseen nodes [5]. In contrast, spectral GNNs rely on the graph Laplacian and Fourier transform for convolution operations, thus learning from global structural patterns. Here we review representative spectral methods. ChebyNet introduces Chebyshev polynomials to construct efficient spectral filters, thereby improving scalability [39]. BernNet learns spectral filters through Bernstein polynomial approximations, which provide flexibility in filter design and help overcome the limitations of predefined or unconstrained filters [40]. JacobiConv increases expressive power by employing an orthogonal Jacobi basis, yielding superior performance over other spectral GNNs across both synthetic and real-world datasets [38]. In brief, spectral GNNs derive filtering expressions that modulate both the low-frequency and high-frequency components of the original graph signals. This process aligns with evolutionary dynamics: low-frequency components represent the stable core of the graph signal, while high-frequency components capture variations among nodes, thereby capturing finer details.
@@
3 Evolution via Spectral Graph Representations
3.1 Preliminary
Notations. Given a set of nodes $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, an undirected graph can be represented in matrix form as $\mathcal{G}=(\mathbf{X},\mathbf{A})$, where the adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ encodes the pairwise interactions between nodes.The degree matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$, and the normalized Laplacian matrix is presented as $\widehat{\mathbf{L}}\ =\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. This matrix can also be expanded in the spectral form as $\widehat{\mathbf{L}}\ =\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}$ contains orthonormal eigenvectors and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$ is a diagonal matrix of the corresponding eigenvalues, which satisfy $\lambda_{k}\in[0,2]$. The graph Fourier transform is defined as $\widetilde{\mathbf{X}}=\mathbf{U}^{\top}\mathbf{X}$, with inverse transform $\mathbf{X}=\mathbf{U}\widetilde{\mathbf{X}}$. The component at frequency $\lambda_{k}$ is given by $\widetilde{\mathbf{X}}_{\lambda_{k}}=\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}$, where $\mathbf{U}_{k}$ is the eigenvector corresponding to $\lambda_{k}$. The spectral filter is defined as $g(\mathbf{\Lambda})=\mathrm{diag}\bigl{(}g(\lambda_{1}),\dots,g(\lambda_{N})\bigr{)}$ and is generally parameterized by a $k$th-order polynomial, $g(\lambda)=\sum_{k=0}^{K}\alpha_{k}\lambda^{k}$, which lifts to the vertex domain through $g(\widehat{\mathbf{L}})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}$.
Problem formulation. In the proposed GNE framework, the evolution of the population can be regarded as a frequency filtering process. By controlling the frequency components, the balance between exploration and exploitation is achieved. Specifically, the GNE population-update procedure is based on a polynomial filter and can be defined as follows:
$g(\widehat{\mathbf{L}})\,\mathbf{X}=\mathbf{U}\,g(\mathbf{\Lambda})\,\mathbf{U}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\widetilde{\mathbf{X}}_{\lambda_{k}},$ (1)
@@ 3.1 Preliminary
3.2 The proposed GNE model
Existing representative evolutionary algorithms, such as DE[14], typically operate via three key steps: mutation, crossover, and selection. During the mutation step, a mutant vector $\bm{v}_{i}$ is generated for each individual $\bm{x}_{i}$ in the population:
$\bm{v}_{i}=\bm{x}_{r1}+\alpha_{s}\cdot(\bm{x}_{r2}-\bm{x}_{r3}),$ (2)
@@ 3.2 The proposed GNE model
$\bm{u}_{i,j}=\begin{cases}\bm{v}_{i,j}&\text{if }rand(0,1)\leq\text{p}_{cr}\text{ or }j=j_{r}\\ \bm{x}_{i,j}&\text{otherwise},\end{cases}$ (3)
@@
$\mathbf{x}_{i}^{\prime}=\begin{cases}\mathbf{u}_{i}&\text{if }f(\mathbf{u}_{i})\leq f(\mathbf{x}_{i})\\ \mathbf{x}_{i}&\text{otherwise},\end{cases}$ (4)
@@
$\mathbf{A}_{ij}=\frac{\mathbf{Z}_{i:}\mathbf{Z}_{j:}^{\top}}{\|\mathbf{Z}_{i:}\|\|\mathbf{Z}_{j:}\|}.$ (5)
@@
$\lambda_{k}=\frac{1}{2}\sum_{ij}\mathbf{A}_{ij}\left(\frac{\mathbf{U}_{ki}}{\sqrt{d_{i}}}-\frac{\mathbf{U}_{kj}}{\sqrt{d_{j}}}\right)^{2},$ (6)
@@
1:  Input: Population size $N$, maximum number of iterations $T$, sampling functions $\varphi(\cdot)$ and $\phi(\cdot)$, spectral filter $g(\cdot)$, objective function $f(\cdot)$.
@@
5:  $\mathbf{x}_{\mathrm{best}}\leftarrow\arg\min_{\mathbf{x}_{i}\in\mathbf{X}}f(\mathbf{x}_{i})$.
6:  $f_{\mathrm{best}}\leftarrow f(\mathbf{x}_{\mathrm{best}})$.
@@
14:     Apply encoder $\phi$ to $\mathbf{X}$: $\mathbf{X}\leftarrow\phi(\hat{\mathbf{X}})$.
@@
$\hat{\mathbf{X}}=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\widetilde{\mathbf{X}}_{\lambda_{k}}.$ (7)
where the term $\widetilde{\mathbf{X}}_{\lambda_{k}}$ represents the population’s component corresponding to $\lambda_{k}$ , while $g(\lambda_{k})$ is the weight applied to this component. Increasing the weight of high-frequency components boosts exploration, while increasing the weight of low-frequency components strengthens exploitation. The updated population $\hat{\mathbf{X}}$ is obtained by integrating different frequency components.
GNNs typically employ multilayer perceptron (MLP) encoders $\psi(\cdot)$ and $\phi(\cdot)$ to transform and map features both before and after filtering, thereby further optimizing node representations. Similarly, GNE treats $\psi(\cdot)$ and $\phi(\cdot)$ as sampling functions that guide the population to generate new individuals toward optimizing overall population fitness, this point will be further discussed in Appendix B. To summarize, the overall update formula for GNE can be expressed as follows:
@@
4 Experiments
In this section, we evaluate the performance of GNE, where the filter function is represented using Chebyshev polynomials, against five well-established and widely recognized evolutionary algorithms: GA[12],DE[14], CMA-ES[21], SDAES[22], and RL-SHADE[23]. The comparison is conducted on nine classical benchmark functions to evaluate the convergence and robustness of GNE. Detailed descriptions of these benchmark functions, including their mathematical expressions, bounds, dimensions, and characteristics, are provided in Appendix C.1.
Additionally, we examine GNE’s performance under challenging scenarios, such as when noise is added to the function or when the optimal solution is shifted, to verify its robustness and adaptability in diverse optimization settings. The experiment was conducted using an Intel Core i5-12400 CPU (2.50 GHz), 16 GB of RAM, and the MATLAB 2023a platform for analysis. As an extension, we also analyze the impact of different polynomial bases on GNE’s performance, with the formulations and analysis results of these polynomials presented in Appendix C.2.
@@ 4 Experiments
Without Noise Sphere 2.31E+04±7.10E+03 4.53E-04±1.75E-04 1.22E-05±9.27E-06 9.22E-18±1.43E-17 1.51E-07±2.17E-07 3.07E-20±5.26E-21
@@
With Noise Sphere 2.31E+04±7.44E+03 2.87E-01±4.21E-02 1.12E-01±2.67E-02 3.64E-01±9.54E-02 2.13E-01±7.92E-02 1.11E-04±8.19E-05
@@
4.1 Comparison of GNE and Other Evolutionary Algorithms
@@
4.2 Robustness of GNE in Noisy Function Environments
In real-world scenarios, function evaluations are often affected by measurement errors and uncertainties[41, 42]. In this section, we introduce random noise uniformly distributed in the range [0, 1] to the nine benchmark functions to evaluate the performance of GNE under noisy function conditions. This setup tests the algorithm’s robustness when noise is present in the computation of function values.
To demonstrate the balance between overall performance and computational time in GNE, Figure 4 presents the Friedman ranking[43] and average running times of different algorithms on benchmark functions under both noise-free and noisy conditions. GA and DE, two classic and simple algorithms, exhibit the shortest running times in both conditions but achieve the lowest rankings. CMA-ES and RL-SHADE rank in the middle under both conditions but require the longest and second-longest running times, respectively.SDAES achieves the shortest running time in the noise-free condition and ranks second, slightly behind GNE. However, in the presence of noise, its optimization efficiency significantly deteriorates, dropping to the second-to-last position, even performing worse than the simple algorithm DE. In contrast, GNE achieves the highest ranking under both noise-free and noisy conditions while using a moderate amount of computational time. Additionally, GNE demonstrates strong adaptability to the optimization of functions with added noise, maintaining robust performance across varying conditions.
4.3 Robustness of GNE for Optimal Solution Deviation
A critical problem in benchmarking and analysis of evolutionary computation methods is that many benchmark functions have optima located at the center of the feasible set, which can create biases in the evaluation of evolutionary algorithms[44].The center-bias in some methods allows them to easily find optima, rendering comparisons with other algorithms that lack such biases ineffective. To address this issue, we study whether the proposed GNE algorithm exhibits overfitting to the optimal solution by testing its performance under varying deviations. The optimal solution is intentionally shifted by deviations of -10, -30, -50, -70, -90, +10, +30, +50, +70, and +90. This setup is designed to test GNE’s robustness when the optimal solution deviates from its original position.Table 2 presents the performance comparison of GNE with other evolutionary algorithms (GA, DE, CMA-ES, SDAES, and RL-SHADE) under different optimal solution deviations in the Sphere function. This highlights the robustness and adaptability of GNE in handling dynamic changes in the solution landscape, thanks to the translation invariance property of its adjacency matrix, which captures global correlations by modeling relative relationships within the population without relying on the boundaries of the solution space, as discussed in Appendix D. Furthermore, GNE guides population updates using global-scale information, mitigating the risk of overfitting caused by the influence of a few individuals on population update.

### v4

@@ Introduction
Graph Neural Networks (GNNs) have been widely applied in fields such as social network analysis, recommendation systems, and protein structure prediction because of their capacity to effectively model complex relational structures using graph representations(Kipf and Welling 2016; Veličković et al. 2017; Xu et al. 2018; Wang et al. 2019; Hamilton, Ying, and Leskovec 2017). In GNNs, dependencies between nodes are captured by propagating and aggregating information across the network, thereby enabling adaptive learning and optimization based on node interactions. Similarly, biological evolution involves genetic interactions characterized by variation, recombination, and natural selection, which are processes that collectively facilitate the evolutionary optimization of genotypes toward greater fitness(Wright et al. 1932; Darwin 2023; Brady 1985). Given these parallels, we propose that GNNs can be understood through an evolutionary perspective, where the propagation of information among nodes mirrors the diversity-exploring mechanisms of crossover and mutation, as well as the adaptive optimization mechanism of directional selection in evolution.
@@
Evolutionary adaptation thrives on two complementary strategies: exploration through genetic diversity (e.g., avian wing polymorphism spanning short, medium, and long morphotypes) and exploitation via trait optimization(Eiben and Schippers 1998; Črepinšek, Liu, and Mernik 2013; Yu, Zhang, and Sun 2024).Consider three avian archetypes: short-winged penguins excel in aquatic foraging, medium-winged gulls balance aerial maneuverability with diving efficiency, and long-winged albatrosses dominate transoceanic flight—each adaptation emerging from mutation-driven exploration and sustained through environmental selection. Similarly, giraffes’ neck elongation, refined iteratively over generations, exemplifies exploitation’s role in amplifying fitness-critical traits. Evolutionary algorithms (EAs) formalize this duality: crossover/mutation operators probe global solution spaces (exploration), while selection pressures refine local optima (exploitation)(Holland 1992; Chao et al. 2024; Storn and Price 1997).
@@
Conventional EAs fail to fully exploit the global relationships among individuals in a population, relying on limited information from a few individuals for updating a single solution(Gorges-Schleuter 1991; Črepinšek, Liu, and Mernik 2013; Opara and Arabas 2019), thus hindering comprehensive learning. Additionally, existing EAs optimize functions in a black-box manner, with parameter selection relying on extensive experimental results. Inappropriate parameter choices can easily disrupt the balance between exploration and exploitation, potentially leading to suboptimal results (Ojha, Timmis, and Nicosia 2022; Tanabe and Fukunaga 2019; Karafotias, Hoogendoorn, and Eiben 2014).To address these limitations, we reinterpret evolutionary algorithms through a GNN lens, proposing Graph Neural Evolution (GNE). In GNE, each candidate solution in population is represented as a node in a graph, with the corresponding value in the adjacency matrix between two nodes reflecting the similarity of them. This approach fully leverages global correlations among individuals within the population, overcoming the constraints of local updates. These global correlations resemble biological reproductive isolation, where genetically similar individuals preferentially exchange traits(Dobzhansky 1982). Moreover, we compute the normalized Laplacian matrix from the population’s adjacency matrix and decompose it into frequency components through its eigenvalues and eigenvectors in the spectral domain. Similar to spectral GNNs, GNE balances high-frequency exploration (diverse traits) and low-frequency exploitation (consistent traits) via filter functions. Notably, this explicit spectral aggregation strategy can control global-scale features through frequency components, making the exploration–exploitation control both interpretable and effective.
@@
This spectral perspective provides a unified mathematical framework: the balance between exploration and exploitation in evolutionary algorithms is equivalent to designing an appropriate filter that assigns suitable weights to the different frequency components of the population, where strengthening the high-frequency components corresponds to reinforcing exploration behavior, and strengthening the low-frequency components corresponds to reinforcing exploitation behavior. By formalizing this duality, GNE provides interpretable control over the exploration-exploitation balance while benefiting from global correlations. Through rigorous evaluations on nine benchmark functions (e.g., Sphere, Rastrigin, and Rosenbrock), GNE consistently outperforms classical algorithms, including Genetic Algorithm (GA) (Holland 1992), Differential Evolution (DE) (Storn and Price 1997), Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen, Müller, and Koumoutsakos 2003), and advanced algorithms such as Evolution Strategy Based on Search Direction Adaptation (SDAES) (He et al. 2019) and Learning Adaptive Differential Evolution by Natural Evolution Strategies (RL-SHADE) (Zhang et al. 2022) across scenarios with original, noise-corrupted, and deviated-optimum conditions. GNE leads other algorithms by multiple orders of magnitude in solution quality (e.g., 3.07e-20 vs. 1.51e-07 mean on Sphere). This framework is broadly applicable to complex optimization problems in the real world, such as engineering design, and optimization of parameters and architecture of neural networks(Cai, Gao, and Li 2019; Zhou et al. 2021; Li et al. 2023).To summarize, our contributions are threefold:
@@
Evolutionary Algorithms. EAs are a family of metaheuristic optimization techniques inspired by natural evolutionary processes, incorporating operators such as natural selection and mutation. GA is one of the earliest and most representative EAs, and is particularly effective in solving discrete and multimodal optimization problems(Holland 1992). In recent years, DE(Storn and Price 1997; Opara and Arabas 2019; Pant et al. 2020) and Evolution Strategy (ES)(Beyer and Schwefel 2002; Lange, Tian, and Tang 2024; Beyer and Arnold 2001) have emerged as prominent research domains. DE generates candidate solutions through differential mutation and selects the fittest among them. Numerous variants have been developed, including Adaptive Differential Evolution with Optional External Archive (JADE) (Zhang and Sanderson 2009), which incorporates external archives and adaptive parameter updates; Success-History Based Parameter Adaptation for Differential Evolution (SHADE) (Tanabe and Fukunaga 2013), which utilizes historically successful parameter settings to refine control parameters; and RL-SHADE (Zhang et al. 2022), which further enhances DE performance by integrating natural evolution strategies for more effective parameter optimization. Among ES methods, Covariance Matrix Adaptation Evolution Strategy (CMA-ES) is one of the most widely recognized algorithms (Hansen, Müller, and Koumoutsakos 2003), adjusting the covariance matrix to effectively explore the solution space. Several variants have been developed, including OpenES for large-scale parallel computation (Salimans et al. 2017); Parameter-Exploring Policy Gradients (PEPG) for estimating gradients in the parameter space (Sehnke et al. 2010);Mixture Model based Evolution Strategy (MMES) (He, Zheng, and Zhou 2020), which accelerates CMA-ES in large-scale optimization tasks; and SDAES (He et al. 2019), which reduces both time and space complexity in high-dimensional settings. From GA to DE, ES, and their variants, the design of evolutionary algorithms has consistently centered on balancing exploration and exploitation. Exploration introduces diversity to expand the search space and prevent premature convergence to local optima, while exploitation refines promising solutions to gradually approach the global optimum(Eiben and Schippers 1998).
@@
Graph Neural Networks. GNNs process graph-structured data from two principal perspectives: spatial and spectral(Chen et al. 2023; Muhammet et al. 2020; Wang and Zhang 2022). Spatial GNNs are more widely known, with the Graph Convolutional Network (GCN) (Kipf and Welling 2016) being a representative example, which aggregates information from neighboring nodes to learn local structures. The Graph Attention Network (GAT) (Veličković et al. 2017) further enhances GCNs by incorporating self-attention mechanisms, while GraphSAGE generates node embeddings by sampling and aggregating features from local neighborhoods, thereby enabling generalization to previously unseen nodes (Hamilton, Ying, and Leskovec 2017). In contrast, spectral GNNs rely on the graph Laplacian and Fourier transform for convolution operations, thus learning from global structural patterns. Here we review representative spectral methods. ChebyNet introduces Chebyshev polynomials to construct efficient spectral filters, thereby improving scalability (Defferrard, Bresson, and Vandergheynst 2016). BernNet learns spectral filters through Bernstein polynomial approximations, which provide flexibility in filter design and help overcome the limitations of predefined or unconstrained filters (He et al. 2021). JacobiConv increases expressive power by employing an orthogonal Jacobi basis, yielding superior performance over other spectral GNNs across both synthetic and real-world datasets (Wang and Zhang 2022). In brief, spectral GNNs derive filtering expressions that modulate both the low-frequency and high-frequency components of the original graph signals. This process aligns with evolutionary dynamics: low-frequency components represent the stable core of the graph signal, while high-frequency components capture variations among nodes, thereby capturing finer details.
@@
Evolution via Spectral Graph Representations
Preliminary
Notations. Given a set of nodes $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, an undirected graph can be represented in matrix form as $\mathcal{G}=(\mathbf{X},\mathbf{A})$, where the adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ encodes the pairwise interactions between nodes.The degree matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$, and the normalized Laplacian matrix is presented as $\widehat{\mathbf{L}}\ =\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. This matrix can also be expanded in the spectral form as $\widehat{\mathbf{L}}\ =\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}$ contains orthonormal eigenvectors and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$ is a diagonal matrix of the corresponding eigenvalues, which satisfy $\lambda_{k}\in[0,2]$. The graph Fourier transform is defined as $\widetilde{\mathbf{X}}=\mathbf{U}^{\top}\mathbf{X}$, with inverse transform $\mathbf{X}=\mathbf{U}\widetilde{\mathbf{X}}$. The component at frequency $\lambda_{k}$ is given by $\widetilde{\mathbf{X}}_{\lambda_{k}}=\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}$, where $\mathbf{U}_{k}$ is the eigenvector corresponding to $\lambda_{k}$. The spectral filter is defined as $g(\mathbf{\Lambda})=\mathrm{diag}\bigl{(}g(\lambda_{1}),\dots,g(\lambda_{N})\bigr{)}$ and is generally parameterized by a $k$th-order polynomial, $g(\lambda)=\sum_{k=0}^{K}\alpha_{k}\lambda^{k}$, which lifts to the vertex domain through $g(\widehat{\mathbf{L}})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}$. Problem formulation. In the proposed GNE framework, the evolution of the population can be regarded as a frequency filtering process. By controlling the frequency components, the balance between exploration and exploitation is achieved. Specifically, the GNE population-update procedure is based on a polynomial filter and can be defined as follows:
$\begin{split}g(\widehat{\mathbf{L}})\,\mathbf{X}=\mathbf{U}\,g(\mathbf{\Lambda})\,\mathbf{U}^{\top}\mathbf{X}=\\ \sum_{k=1}^{n}g(\lambda_{k})\,\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\widetilde{\mathbf{X}}_{\lambda_{k}}\end{split}$ (1)
@@
The proposed GNE model
Existing representative evolutionary algorithms, such as DE(Storn and Price 1997), typically operate via three key steps: mutation, crossover, and selection. During the mutation step, a mutant vector $\bm{v}_{i}$ is generated for each individual $\bm{x}_{i}$ in the population:
$\bm{v}_{i}=\bm{x}_{r1}+\alpha_{s}\cdot(\bm{x}_{r2}-\bm{x}_{r3})$ (2)
@@
$\bm{u}_{i,j}=\begin{cases}\bm{v}_{i,j}&\text{if }rand(0,1)\leq\text{p}_{cr}\text{ or }j=j_{r}~,\\ \bm{x}_{i,j}&\text{otherwise}\end{cases}$ (3)
@@
$\mathbf{x}_{i}^{\prime}=\begin{cases}\mathbf{u}_{i}&\text{if }f(\mathbf{u}_{i})\leq f(\mathbf{x}_{i})~,\\ \mathbf{x}_{i}&\text{otherwise}\end{cases}$ (4)
@@
$\mathbf{A}_{ij}=\frac{\mathbf{Z}_{i:}\mathbf{Z}_{j:}^{\top}}{\|\mathbf{Z}_{i:}\|\|\mathbf{Z}_{j:}\|}$ (5)
@@
$\lambda_{k}=\frac{1}{2}\sum_{ij}\mathbf{A}_{ij}\left(\frac{\mathbf{U}_{ki}}{\sqrt{d_{i}}}-\frac{\mathbf{U}_{kj}}{\sqrt{d_{j}}}\right)^{2}$ (6)
@@
1:Input: Population size NN, maximum number of iterations $T$, sampling functions $\varphi(\cdot)$ and $\phi(\cdot)$, spectral filter $g(\cdot)$, objective function $f(\cdot)$.
@@
[…]
@@
[…]
@@
$\hat{\mathbf{X}}=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}=\sum_{k=1}^{n}g(\lambda_{k})\,\widetilde{\mathbf{X}}_{\lambda_{k}}$ (7)
where the term $\widetilde{\mathbf{X}}_{\lambda_{k}}$ represents the population’s component corresponding to $\lambda_{k}$ , while $g(\lambda_{k})$ is the weight applied to this component. Increasing the weight of high-frequency components boosts exploration, while increasing the weight of low-frequency components strengthens exploitation. The updated population $\hat{\mathbf{X}}$ is obtained by integrating different frequency components. GNNs typically employ multilayer perceptron (MLP) encoders $\psi(\cdot)$ and $\phi(\cdot)$ to transform and map features both before and after filtering, thereby further optimizing node representations. Similarly, GNE treats $\psi(\cdot)$ and $\phi(\cdot)$ as sampling functions that guide the population to generate new individuals toward optimizing overall population fitness. Here, $\psi(\cdot)$ is an identity transformation, while $\phi(\cdot)$ is a Gaussian distribution centered at the current best solution $\mathbf{x}_{\text{best}}$ with variance $0.0005\cdot\exp\left(-16\cdot\frac{t-T/2}{T/2}\right)$, where $t$ is the current iteration and $T$ is the maximum number of iterations.To summarize, the overall update formula for GNE can be expressed as follows:
@@
Without Noise
Sphere 2.31E+04±7.10E+03 4.53E-04±1.75E-04 1.22E-05±9.27E-06 9.22E-18±1.43E-17 1.51E-07±2.17E-07 3.07E-20±5.26E-21
@@
With Noise
Sphere 2.31E+04±7.44E+03 2.87E-01±4.21E-02 1.12E-01±2.67E-02 3.64E-01±9.54E-02 2.13E-01±7.92E-02 1.11E-04±8.19E-05
@@
Experiments
In this section, we evaluate the performance of GNE, where the filter function is represented using Chebyshev polynomials, against five well-established and widely recognized evolutionary algorithms: GA(Holland 1992),DE(Storn and Price 1997), CMA-ES(Hansen, Müller, and Koumoutsakos 2003), SDAES(He et al. 2019), and RL-SHADE(Zhang et al. 2022). The comparison is conducted on nine classical benchmark functions to evaluate the convergence and robustness of GNE. Detailed descriptions of these benchmark functions, including their mathematical expressions, bounds, dimensions, and characteristics.
Additionally, we examine GNE’s performance under challenging scenarios, such as when noise is added to the function or when the optimal solution is shifted, to verify its robustness and adaptability in diverse optimization settings. The experiment was conducted using an Intel Core i5-12400 CPU (2.50 GHz), 16 GB of RAM, and the MATLAB 2023a platform for analysis.
Comparison of GNE and Other Evolutionary Algorithms